miggy14671 miggy14671
  • 01-02-2022
  • Mathematics
contestada

A fair die is tossed twice. Find the probability that a 3 is rolled at least once in the two tosses.

Respuesta :

jcherry99
jcherry99 jcherry99
  • 06-02-2022

Answer:

Step-by-step explanation:

Here are the "winning" rolls

1/6 * 5/6 = 5/36

5/6 * 1/6 = 5/36

1/6 * 1/6 = 1/36

There are three tosses out of 36 that are successful.

11/36 = 0.306

Answer Link

Otras preguntas

Which is the same as dividing 35 by 6
What is the long term effect of tariffs and other trade barriers ?
Which organelle is most prominent in the cells and makes protein synthesis?
Suppose the diameter of a circle is 4. What is its radius? and please explain
PLEASE MATH HELP WILL GIVE BRAINLIEST Solve for x. Enter your answer in the box.
Cystic fibrosis (cf) is caused by a recessive allele. a couple who do not have cf have a child that has cf. what is the probability that if they have another ch
What is the definition of point of tangency? in Geometry
“Na2SO3” represents sodium sulfite. What does the 4 mean in the formula 4Na2SO3? A. There are four atoms of sodium. B. There are four atoms of sodium sulfite. C
If the president wants to engage in military action A the president can send troops into action, but Congress must declare war B the Senate can send troops in
How much greater is 300 than 3